# Control of spatiotemporal chaos by stochastic resetting

### v1

@@ Abstract
Discussions and outlook.– We have shown that stochastic resetting provides a powerful mechanism for mitigating spatiotemporal chaos and can even induce a dynamical phase transition to a non-chaotic regime, where the spreading of information is brought to a complete halt. This can be put in perspective with the quantum many-body localized regimes where, although slower than ballistic, the spreading of information is thought to survive in the absence of transport. More generally, the insights gained from these classical results will pave the path to understanding ergodicity and chaos in many-body quantum systems subject to resetting [76, 77, 78, 79]. While we focused on discrete nonlinear maps, our results readily adapt to continuous-time dynamics. Using the dictionary $\lambda\to\lambda\,{\rm d}t$ and $r\to r\,{\rm d}t$, the renormalized Lyapunov in Eq. (9) becomes $\tilde{\lambda}=\lambda-r$, the butterfly velocity in Eq. (17) becomes $\tilde{v}_{\rm B}=\lambda^{-1}(r)$, and the expression of $\tilde{\lambda}_{\rm B}$ in (19) remains valid, see Ref. [61] for more details. The versatility of these results suggests the possibility of developing a geometrical optics theory to investigate phenomena such as the refraction of “information light rays” on spatial domains subject to resetting.
Acknowledgements.– M.K. acknowledges support from the Department of Atomic Energy, Government of India, under project No. RTI4001. M.K. also thanks the VAJRA faculty scheme (No. VJR/2019/000079) from the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India. We gratefully acknowledge support from the International Research Project (IRP) titled ‘Classical and Quantum Dynamics in Out of Equilibrium Systems’ (DynoutSys), funded by CNRS, France. M.K. thanks the Institute of Physics at EPFL in Lausanne for their hospitality. C.A. thanks the International Centre for Theoretical Sciences (ICTS) in Bangalore for their hospitality. This research was supported in part by the ICTS and CEFIPRA for participating in the “Indo-French Workshop on Classical and quantum dynamics in out of equilibrium systems” (ICTS/IFWCQM2024/12).
Supplementary Material
I Discrete-time maps subject to stochastic resetting
I.1 Stationary distribution of the deterministic map
@@
$p_{\rm st}(x)=\lim\limits_{N\to\infty}\frac{1}{N+1}\sum_{n=0}^{N}\delta(x-x_{n})$ (S2)
@@
I.2 Stationary distribution with stochastic resetting
In the presence of a finite resetting rate ($10<r<$ 1), the dynamics are generated by the following stochastic map
$x_{n}\mapsto x_{n+1}=\left\{\begin{array}[]{ll}f(x_{n})&\text{with probability }1-r\\ x_{0}&\text{with probability }r\,.\end{array}\right.$ (S5)
The corresponding stationary distribution, if it exists, is given by the time-averaged distribution $\frac{1}{m+1}\sum_{n=0}^{m}\delta(x-x_{n})$ truncated at the time $m$ where resetting occurs, and averaged over all possible resetting times:
$\tilde{p}_{\rm st}(x)=\sum_{m=0}^{\infty}P_{m}\frac{1}{m+1}\sum_{n=0}^{m}\delta(x-x_{n})$ (S6)
with $P_{m}:=r(1-r)^{m}$ the probability of a reset immediately after $m$ time steps without reset. Standard algebra yields
$\tilde{p}_{\rm st}(x)=\sum_{n=0}^{\infty}\tilde{g}_{n}\,\delta(x-x_{n})\,,$ (S7)
with $\tilde{g}_{n}:=r(1-r)^{n}\,\Phi(1-r,1,1+n)$ where $\Phi$ is the Hurwitz–Lerch transcendent. $g_{n}$ is monotonously decreasing with $n$ and $\tilde{g}_{n}\sim\exp(-r\,n)/n$ at large $n$. One may check the normalization
$\int{\rm d}x\,\tilde{p}_{\rm st}(x)=\sum_{n=0}^{\infty}\tilde{g}_{n}=1\,.$ (S8)
Resetting is naturally expected to enhance the overall time spent by the system in the initial state, $x_{0}$, and its first subsequent images under the map $f$. This is confirmed by examining Eq. (S7), which, unlike the deterministic case in Eq. (S2), lacks invariance under arbitrary permutations of the $x_{n}$’s. Notably, the initial condition exhibits the highest weight to the stationary distribution $\tilde{p}_{\rm st}(x)$, given by $\tilde{g}_{0}=\frac{r}{1-r}\log(1/r)$. This demonstrates the dependence of the stationary distribution $\tilde{p}_{\rm st}(x)$ on the choice of the initial condition.
@@ I.2 Stationary distribution with stochastic resetting
This remarkable result states that, under such an averaging procedure of the initial conditions, the statics of the map are not altered by stochastic resetting. For instance, this implies that the celebrated bifurcation diagram of the logistic map is not altered by stochastic resetting. However, in this work, we do not perform this additional averaging over the initial conditions. Instead, we focus on the case of a single, arbitrary, initial condition $x_{0}$. We have carefully checked that our results do not depend on the choice of $x_{0}$.
I.3 Ergodicity with stochastic resetting
@@
Ergodicity of the deterministic map, i.e. $p_{\rm st}(x)>0$ for arbitrary generic values of $x$, implies that for a given small $\epsilon>0$, and a large enough time $N$, the condition $|x-x_{n}|<\epsilon$ is met by a finite set $\Lambda_{\epsilon,N}(x)$ of $n$’s along the trajectory before time $N$ is reached. The size of this set is on the order of $\mathcal{O}(\epsilon\times N)$. In practice, we question whether the stationary distribution in the presence of stochastic resetting $\tilde{p}_{\rm st}(x)>0$ for arbitrary generic values of $x$. Integrating Eq. (S7) on a small interval of length $2\epsilon$ around $x$, we get
$\int_{x-\epsilon}^{x+\epsilon}\!\!\!\!{\rm d}y\,\tilde{p}_{\rm st}(y)=\sum_{n\in\Lambda_{\epsilon}(y)}\tilde{g}_{n}\,,$ (S9)
@@
$\tilde{p}_{\rm st}(x)=\frac{1}{2\epsilon}\sum_{n\in\Lambda_{\epsilon}(x)}\tilde{g}_{n}\,.$ (S10)
@@
$\tilde{p}_{\rm st}(x)>\frac{1}{2\epsilon}\tilde{g}_{n_{\rm min}}\,,$ (S11)
where $n_{\rm min}$ is chosen to be the smallest element of $\Lambda_{\epsilon}(x)$, corresponding to the largest weight $\tilde{g}_{n}$ of the sum in Eq. (S10). We now use the chaotic character of the deterministic dynamics to get an estimate for $n_{\rm min}$: the typical distance between two trajectories originating from neighboring initial conditions at a small enough distance $\epsilon$ grows exponentially
$\frac{|\delta x_{n}|}{\epsilon}\sim{\rm e}^{\lambda n}\text{ at large enough $n$,}$ (S12)
@@
$n_{\rm min}\sim\frac{1}{\lambda}\log\frac{1}{\epsilon}\,.$ (S13)
Using the asymptotic expression of $\tilde{g}_{n}$, this yields the bound
$\tilde{p}_{\rm st}(x)\gtrsim\frac{\lambda}{2\log 1/\epsilon}\,\left(\frac{1}{\epsilon}\right)^{\frac{\log(1-r)}{\lambda}+1}\,.$ (S14)
@@
II Deterministic coupled logistic map
@@
$F_{i}(x_{n})=f(x_{n,\,i})+\frac{c}{2}\left[f(x_{n,\,{i-1}})-2f(x_{n,\,i})+f(x_{n,\,{i+1}})\right]\,,$ (S16)
together with periodic boundary conditions. Here, $f(x)=\alpha\,x(1-x)$ is the logistic function with the parameter $\alpha\in[0,4]$ and $10<c<$ 1 is a diffusion parameter. This ensures the multidimensional interval $[0,1]^{L}$ is stable under the coupled logistic map. In practice, we work with $\alpha=4$ and $c=0.1$ for which we have carefully checked that the deterministic dynamics reach a stationary state that is both uniform in space and time. This is illustrated in Fig. S1 where the local marginal of the stationary distribution is plotted at sites $i=1$ and $i=(L-1)/2$. For comparison, we also plot $p_{\rm st}(x)=1/(\pi\sqrt{x(1-x)})$ which is the stationary distribution of the logistic map (at $c=0$).
Figure S1: Coupled logistic map with $\alpha=4$ and $c=0.1$ (no resetting, $r=0$): the local marginal of the stationary distribution is identical at all sites, here $i=1$ and $i=(L-1)/2$. The stationary distribution of the corresponding logistic map ($c=0$), $p_{\rm st}(x)=1/(\pi\sqrt{x(1-x)})$, is plotted for comparison. Lattice size $L=701$.
@@
$D_{n,\,ij}=\left|\frac{\delta x_{n,\,i}}{\delta x_{0,\,j}}\right|\,.$ (S17)
@@
$D_{n,\,ij}=|\left\{F^{\prime}(x_{0})\times F^{\prime}(x_{1})\times\ldots\times F^{\prime}(x_{n-1})\right\}_{ij}|\,,$ (S18)
@@
$F_{ij}^{\prime}(x_{n}):=\frac{\partial F_{i}(x_{n})}{\partial x_{n,\,j}}\,,$ (S19)
@@
$F_{ij}^{\prime}(x)=(1-c)\frac{\partial f(x_{i})}{\partial x_{j}}\delta_{ij}+\frac{c}{2}\!\left[\frac{\partial f(x_{i})}{\partial x_{j+1}}\delta_{i,j+1}+\frac{\partial f(x_{i})}{\partial x_{j-1}}\delta_{i,j-1}\right]\,,$ (S20)
@@
III Continuous-time dynamics subject to stochastic resetting
@@
III.1 Zero-dimensional systems (no space)
@@
$\begin{array}[]{rl}n&\longleftrightarrow t/{\rm d}t\\ x_{n}&\longleftrightarrow x(t)\\ x_{n+1}-x_{n}&\longleftrightarrow\partial_{t}x(t)\end{array}\qquad\begin{array}[]{rl}\lambda&\longleftrightarrow\lambda\,{\rm d}t\\ r&\longleftrightarrow r\,{\rm d}t\\ (1-r)^{n}&\longleftrightarrow{\rm e}^{-rt}\end{array}$ (S27)
where $\lambda$ and $r$ are now rates instead of dimensionless exponent and probability, respectively. In particular, the resetting rate $r$ now takes values in $[0,\infty)$.
III.1.1 Deterministic dynamics & Lyapunov rate
Let us consider deterministic dynamics governed by the following first-order PDE
$\partial_{t}x(t)=f(x(t))\,,$ (S28)
where $f(x)$ is a non-linear function of $x$ which is assumed to generate ergodic and chaotic dynamics.
@@ III.1.1 Deterministic dynamics & Lyapunov rate
$d(t):=\left|\frac{\delta x(t)}{\delta x(0)}\right|\,.$ (S29)
[…]
@@
$\partial_{t}\delta x(t)=\delta x(t)f^{\prime}(x(t))\,,$ (S30)
@@
$\lim\limits_{t\to\infty}\frac{1}{t}\int_{0}^{t}{\rm d}\tau\,|f^{\prime}(x(\tau))|=\int{\rm d}x\,p_{\rm st}(x)|f^{\prime}(x)|=:\lambda>0\,.$ (S32)
which provides a static definition of the Lyapunov rate. It connects to the traditional dynamical definition of the Lyapunov via the asymptotic expression at large times $t\gg 1/\lambda$
$d(t)\simeq{\rm e}^{\lambda t}\,.$ (S33)
III.1.2 Dynamics subject to stochastic resetting & Lyapunov rate

### v2

@@ III Coupled map lattices
IV Discussions and outlook
We have shown that stochastic resetting provides a powerful mechanism for mitigating spatiotemporal chaos and can even induce a dynamical phase transition to a non-chaotic regime, where the spreading of information is brought to a complete halt. This can be put in perspective with the quantum many-body localized regimes where, although slower than ballistic, the spreading of information is thought to survive in the absence of transport. More generally, the insights gained from these classical results will pave the path to understanding ergodicity and chaos in many-body quantum systems subject to resetting [83, 84, 85, 86]. While we focused on discrete nonlinear maps, our results readily adapt to continuous-time dynamics. Using the dictionary $\lambda\to\lambda\,{\rm d}t$ and $r\to r\,{\rm d}t$, the renormalized Lyapunov in Eq. (9) becomes $\tilde{\lambda}=\lambda-r$, the butterfly velocity in Eq. (17) becomes $\tilde{v}_{\rm B}=\lambda^{-1}(r)$, and the expression of $\tilde{\lambda}_{\rm B}$ in (19) remains valid, see Appendix C for more details. The versatility of these results suggests the possibility of developing a geometrical optics theory to investigate phenomena such as the refraction of “information light rays” on spatial domains subject to resetting.
Acknowledgements
M.K. acknowledges support from the Department of Atomic Energy, Government of India, under project No. RTI4001. M.K. also thanks the VAJRA faculty scheme (No. VJR/2019/000079) from the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India. We gratefully acknowledge support from the International Research Project (IRP) titled ‘Classical and Quantum Dynamics in Out of Equilibrium Systems’ (DynoutSys), funded by CNRS, France. M.K. thanks the Institute of Physics at EPFL in Lausanne for their hospitality. C.A. thanks the International Centre for Theoretical Sciences (ICTS) in Bangalore for their hospitality. This research was supported in part by the ICTS and CEFIPRA for participating in the “Indo-French Workshop on Classical and quantum dynamics in out of equilibrium systems” (ICTS/IFWCQM2024/12).
Appendix A Discrete-time maps subject to stochastic resetting
A.1 Stationary distribution of the deterministic map
@@
$p_{\rm st}(x)=\lim\limits_{N\to\infty}\frac{1}{N+1}\sum_{n=0}^{N}\delta(x-x_{n})$ (22)
@@
A.2 Stationary distribution with stochastic resetting
In the presence of a finite resetting probability ($0<r<1$), the dynamics are generated by the following stochastic map
$x_{n}\mapsto x_{n+1}=\left\{\begin{array}[]{ll}f(x_{n})&\text{with probability }1-r\\ x_{0}&\text{with probability }r\,.\end{array}\right$. (25)
The corresponding stationary distribution, if it exists, is given by renewal equation
$\tilde{p}_{\rm st}(x)=\sum_{m=0}^{\infty}\tilde{g}_{m}\,\delta(x-f^{m}(x_{0}))\,$, (26)
where $\tilde{g}_{m}:=r(1-r)^{m}$ is the probability that a reset occurred exactly $m$ time steps ago, and that no reset has occurred in the subsequent $m$ steps. One may check the normalization
$\int{\rm d}x\,\tilde{p}_{\rm st}(x)=\sum_{m=0}^{\infty}\tilde{g}_{m}=1\,.$ (27)
Resetting is naturally expected to enhance the overall time spent by the system in the initial state, $x_{0}$, and its first subsequent images under the map $f$. This is confirmed by examining Eq. (26), which, unlike the deterministic case in Eq. (22), lacks invariance under arbitrary permutations of the $x_{n}$’s. Notably, the initial condition exhibits the highest weight to the stationary distribution $\tilde{p}_{\rm st}(x)$, given by $\tilde{g}_{0}=r$. This demonstrates the dependence of the stationary distribution $\tilde{p}_{\rm st}(x)$ on the choice of the initial condition.
@@ A.2 Stationary distribution with stochastic resetting
This result states that, under such an averaging procedure of the initial conditions, the statics of the map are not altered by stochastic resetting. For instance, this implies that the celebrated bifurcation diagram of the logistic map is not altered by stochastic resetting. However, in this work, we do not perform this additional averaging over the initial conditions. Instead, we focus on the case of a single, arbitrary, initial condition $x_{0}$. We have carefully checked that our results do not depend on the choice of $x_{0}$.
A.3 Ergodicity with stochastic resetting
@@
By ergodicity, we mean that for almost all initial conditions $x_{0}$, the subsequent system’s trajectory passes arbitrarily close to any point in phase space, ensuring that the system eventually explores all regions of phase space with a frequency proportional to their volume. Note that this definition does not necessarily require independence (of these frequencies) with respect to initial conditions. More precisely, the ergodicity of the deterministic map, means that almost all $x$ in phase space, a given small $\epsilon>0$, and a large enough time $N$, the condition $|x-x_{n}|<\epsilon$ is met by a finite set $\Lambda_{\epsilon,N}(x)$ of $n$’s along the trajectory before time $N$ is reached. The size of this set is on the order of $\mathcal{O}(\epsilon\times N)$, and ergodicity simply implies the nonvanishing of the stationary distribution, $p_{\rm st}(x)>0$, for almost all $x$. In practice, we question whether the stationary distribution in the presence of stochastic resetting is non-vanishing for arbitrary generic values of $x$, i.e. $\tilde{p}_{\rm st}(x)>0$. Integrating Eq. (26) on a small interval of length $2\epsilon$ around $x$, we get
$\int_{x-\epsilon}^{x+\epsilon}\!\!\!\!{\rm d}y\,\tilde{p}_{\rm st}(y)=\sum_{n\in\Lambda_{\epsilon}(y)}\tilde{g}_{n}\,$, (28)
@@
$\tilde{p}_{\rm st}(x)=\frac{1}{2\epsilon}\sum_{n\in\Lambda_{\epsilon}(x)}\tilde{g}_{n}\,.$ (29)
@@
$\tilde{p}_{\rm st}(x)>\frac{1}{2\epsilon}\tilde{g}_{n_{\rm min}}\,$, (30)
where $n_{\rm min}$ is chosen to be the smallest element of $\Lambda_{\epsilon}(x)$, corresponding to the largest weight $\tilde{g}_{n}$ of the sum in Eq. (29). We now use the chaotic character of the deterministic dynamics to get an estimate for $n_{\rm min}$: the typical distance between two trajectories originating from neighboring initial conditions at a small enough distance $\epsilon$ grows exponentially
$\frac{|\delta x_{n}|}{\epsilon}\sim{\rm e}^{\lambda n}\text{ at large enough $n$,}$ (31)
@@
$n_{\rm min}\sim\frac{1}{\lambda}\log\frac{1}{\epsilon}\,.$ (32)
Using the expression of $\tilde{g}_{n}$, this yields the bound
$\tilde{p}_{\rm st}(x)\gtrsim\frac{r}{2}\,\left(\frac{1}{\epsilon}\right)^{\frac{\log(1-r)}{\lambda}+1}\,.$ (33)
@@
Appendix B Deterministic coupled logistic map
@@
$F_{i}(x_{n})=f(x_{n,\,i})+\frac{c}{2}\left[f(x_{n,\,{i-1}})-2f(x_{n,\,i})+f(x_{n,\,{i+1}})\right]\,$, (35)
together with periodic boundary conditions. Here, $f(x)=\alpha\,x(1-x)$ is the logistic function with the parameter $\alpha\in[0,4]$ and $0<c<1$ is a diffusion parameter (the term in brackets is a discrete version of the one-dimensional Laplacian). This ensures the multidimensional interval $0,1]^{L}$ is stable under the coupled logistic map. In practice, we work with $\alpha=4$ and $c=0.1$ for which we have carefully checked that the deterministic dynamics reach a stationary state that is both uniform in space and time. This is illustrated in Fig. 4 where the local marginal of the stationary distribution is plotted at sites $i=1$ and i=(L1)/2i=(L-1)/2. For comparison, we also plot $p_{\rm st}(x)=1/(\pi\sqrt{x(1-x)})$ which is the stationary distribution of the logistic map (at $c=0$).
Figure 4: Coupled logistic map with $\alpha=4$ and $c=0.1$ (no resetting, $r=0$): the local marginal of the stationary distribution is identical at all sites, here $i=1$ and i=(L1)/2i=(L-1)/2. The stationary distribution of the corresponding logistic map ($c=0$), $p_{\rm st}(x)=1/(\pi\sqrt{x(1-x)})$, is plotted for comparison. Lattice size $L=701$.
@@
$D_{n,\,ij}=\left|\frac{\delta x_{n,\,i}}{\delta x_{0,\,j}}\right|\,.$ (36)
@@
$D_{n,\,ij}=|\left\{F^{\prime}(x_{0})\times F^{\prime}(x_{1})\times\ldots\times F^{\prime}(x_{n-1})\right\}_{ij}|\,$, (37)
@@
$F_{ij}^{\prime}(x_{n}):=\frac{\partial F_{i}(x_{n})}{\partial x_{n,\,j}}\,$, (38)
@@
$F_{ij}^{\prime}(x)=$ $(1-c)\frac{\partial f(x_{i})}{\partial x_{j}}\delta_{ij}$
$+\frac{c}{2}\!\left[\frac{\partial f(x_{i})}{\partial x_{j+1}}\delta_{i,j+1}+\frac{\partial f(x_{i})}{\partial x_{j-1}}\delta_{i,j-1}\right]\,$, (39)
@@
Appendix C Continuous-time dynamics subject to stochastic resetting
@@
C.1 Zero-dimensional systems (no space)
@@
$\begin{array}[]{rl}n&\longleftrightarrow t/{\rm d}t\\ x_{n}&\longleftrightarrow x(t)\\ x_{n+1}-x_{n}&\longleftrightarrow\partial_{t}x(t)\end{array}\qquad\begin{array}[]{rl}\lambda&\longleftrightarrow\lambda\,{\rm d}t\\ r&\longleftrightarrow r\,{\rm d}t\\ (1-r)^{n}&\longleftrightarrow{\rm e}^{-rt}\end{array}$ (46)
where $\lambda$ and $r$ are now rates instead of dimensionless exponent and probability, respectively. In particular, the resetting rate $r$ now takes values in $0,\infty)$.
C.1.1 Deterministic dynamics & Lyapunov rate
Let us consider deterministic dynamics governed by the following first-order PDE at times $t>0$
$\partial_{t}x(t)=f(x(t))\,$, (47)
where $f(x)$ is a non-linear function of $x$ which is assumed to generate ergodic and chaotic dynamics, and the initial condition $x(0)$ at $t=0$.
@@ C.1.1 Deterministic dynamics & Lyapunov rate
$d(t):=\left|\frac{\delta x(t)}{\delta x(0)}\right|\,.$ (48)
To make the connection with the study of quantum chaos, the quantity $d(t)$ can be referred to as an “out-of-time-order correlator” (OTOC). Indeed, in a semi-classical sense, ${\delta x(t)}/{\delta x(0)}$ is to be replaced by the commutator $\hat{x}(t),\hat{p}(0)]/{\rm i}\hbar$ where $\hat{p}$ is the momentum operator conjugated with the position operator $\hat{x}$, $\hat{x},\hat{p}]={\rm i}\hbar$. Hence, $d^{2}(t)$ is to be replaced by the square commutator $\langle[\hat{x}(t),\hat{p}(0)]^{\dagger}[\hat{x}(t),\hat{p}(0)]\rangle/({\rm i}\hbar)^{2}$ where quantum mechanical operators explicitly appear out of time order and $\langle\ldots\rangle$ denotes averaging over an appropriately chosen initial density matrix.
@@
$\partial_{t}\delta x(t)=\delta x(t)f^{\prime}(x(t))\,$, (49)
@@
$\lim\limits_{t\to\infty}\frac{1}{t}\int_{0}^{t}{\rm d}\tau\,|f^{\prime}(x(\tau))|=\int{\rm d}x\,p_{\rm st}(x)|f^{\prime}(x)|=:\lambda>0\,.$ (51)
which provides a static statistical expression of the Lyapunov rate. It connects to the traditional dynamical definition of the Lyapunov via the asymptotic expression at large times $t\gg 1/\lambda$
$d(t)\simeq{\rm e}^{\lambda t}\,.$ (52)
C.1.2 Dynamics subject to stochastic resetting & Lyapunov rate